OpenStudy (anonymous):

How would I solve this?: |-2.8| + |2.8|

jigglypuff314 (jigglypuff314):

anything in the absolute value signs would turn positive so |-2.8| = 2.8 so you can get 2.8 + 2.8 = ?

OpenStudy (anonymous):

ok so no matter if its positive or not it will always be positive if it's in the "| |"

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

so 5.6

jigglypuff314 (jigglypuff314):

correct ^_^

OpenStudy (anonymous):

can you help with one more?

jigglypuff314 (jigglypuff314):

of course :)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

5 - [ -(-2) ]

jigglypuff314 (jigglypuff314):

lol yeah, they were having some copy paste issues with fancy characters... know that negative times negative = positive and positive times negative = negative

OpenStudy (anonymous):

so would the [-(-2) ] be -2?

jigglypuff314 (jigglypuff314):

mmm -(-2) = +2 then -[-(-2)] = -[+2] = -2

OpenStudy (anonymous):

so 5 - (-2)?

jigglypuff314 (jigglypuff314):

not quite the stuff inside the outer parenthesis turns into -(-2) = +2 so you get 5 - (+2) =

OpenStudy (anonymous):

7?

jigglypuff314 (jigglypuff314):

not quite, try again 5 - (+2) = 5 - 2 = ? what's 5 minus 2 ? :3

OpenStudy (anonymous):

ooh duh, I was adding the 5 and 2 xD ok cool thanks soo much :)!

jigglypuff314 (jigglypuff314):

haha Glad I could help :)